OpenStudy (anonymous):

to solve by completing the square what value should you add to the equation? x^2+16x=-8

OpenStudy (anonymous):

56 i believe. rewrite: x^2 + 16x + 8 = 0 16 / 2 =8 8^2 = 64 so you need to add 64, but you already added 8. 64-8 = 56

OpenStudy (anonymous):

ok, thanks :)